# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v3

@@ IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling, the dimerization parameter, the number and length of the chain on the exciton transport dynamics by numerically solving the master equation Eq. (13) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by the cavity for strong exciton-cavity coupling, namely the polaritonic branches forming. In order to let the polaritons act as a dominated role in exciton transport, we set that the Rabi splitting is far stronger than the decay rates of the cavity and the TLSs, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
@@ IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relation between the energy level crossings (anticrossings) dressed by the cavity and the exciton transport dynamics to explain the exciton transport mechanism. Based on the relation and the properties of exciton transport dynamics, the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations are expected to be presented.
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. From the Eq. (23), it can be known that the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ of the outgoing sites, thus only $I_{\rm o}$ is shown in following figures, and $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$. One can see that the effects of the inter-chain coupling and the dimerization parameter on exciton transport are significantly different. The steady-state occupation number $p_{11}$ of site ”11” (The two ”1” denote respectively the 1st row and the 1st column in the array of TLSs shown in Fig. 1. This representation is similar below, unless otherwise noted.), $p_{\rm bulk}$ of bulk sites (The ”bulk” contains the all sites from the 2nd to the $(N-1)$th column), the input and outgoing exciton current $I_{\rm i,o}$, and the exciton transport efficiency $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
@@
For the six configurations, the descending sort of efficiency is successively HO-HE ST, HO ST, HO TT, HO-HE TT, HE ST, HE TT in $\zeta\in[-1,1]$ [Fig. 6(a5)]. For the HO-HE ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess commonly relatively high occupation numbers of exciton, which support exciton transport in multichains. Therefore, the exciton transport efficiency, input and outgoing currents are the largest near $\zeta=0.3$ for the HO-HE ST [Fig. 6(a3,4,5)]. On the contrary, for the HO-HE ST at $\zeta=0$ and $\zeta\rightarrow\pm 1$, although the site 11 possesses more excitations [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [see cyan line in Fig. 6(a2)], which is beneficial to energy storage and can be used to design quantum battery quach2022superabsorption . Except for the HE ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as HO TT and HO-HE TT also do near $\zeta=\pm 0.4$. However, for HO ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the ability both of energy transport and storage are strong. Near $\zeta=\pm 0.7$ for HO ST, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
@@
It has four configurations HE ST, HE TT, HO-HE ST, HO-HE TT containing HE chain. Because of the different spectra between odd and even HE chain, in Fig. 6(b1-5) there are two situations $N=8,~{}9$ about the steady-state dynamics with $\delta$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations HE ST and HO-HE ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the HE TT of $N=9$ is the most beneficial to energy storage because of the high occupation number $p_{\rm bulk}$ of bulk sites [Fig. 6(b2)] as well as the lowest exciton currents and transport efficiency. Particularly, the exciton transport efficiency is high for $\delta<0$, but the exciton currents approach to zero, as shown in Fig. 6(b3-5), thus there is no real exciton transport. The reason is that the site 11 can not be coherently excited by drive, namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the HO-HE ST and HO-HE TT of $N=8,9$ or approaches to zero for the others.
@@
In order to investigate the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ on the exciton transport dynamics in the presence of cavity, we plot the Fig. 7 and 8 corresponding to Fig. 6, in which the occupation number $p_{\rm c}$ of photons is added. We pay closely attention to the change of the steady-state dynamics with $\zeta$ and $\delta$ at the crossings (anticrossings) of nearly-zero-energy modes dressed by cavity.
@@
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system in a cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6) respectively. Particularly, to analyse what role the cavity acts as in exciton transport, the dependence of the occupation number $p_{\rm c}$ of photons on $\zeta$ is shown, and its extreme points are marked by the vertical fine lines in Fig. 7(a5), which correspond to the locations of the crossings and anticrossings in spectra, as well as of the incontinuity and extreme points of the second order partial derivative $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of the photonic Hopfield coefficients, marked by the vertical fine lines in Fig. 2 and 3. In consequence, the second order partial derivative of the photonic Hopfield coefficients can also describe the change of the occupation number of photons. It can be seen that the presence of the cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations HO ST, HE ST, HE TT, and HO-HE ST [Fig. 7(a3,4)]. Comparing with Fig. 6, one can discover that the exciton transport is suppressed by the cavity in the range of $\zeta\neq 0$. For example, near $\zeta=-0.3$ $p_{\rm c}$ is high [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for HO ST [Fig. 7(a3,4)]. On the other hand, there are three configurations HE ST, HE TT, and HO-HE TT for which the cavity induces significant increase of $I_{\rm i,o}$ and $\eta$ near $\zeta=0$. In the situation, the exciton transport is assisted by polaritons, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number $p_{N}$ on the last column sites [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Because of few excitations injected [Fig. 7(a1)] for HO ST, HO TT and few $p_{N}$ [Fig. 7(a4)] for HO-HE ST near $\zeta=0$, the cavity has no effect on exciton transport. What’s more, it can be seen obviously that the peaks of $p_{\rm c}$ and of $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ the multichains do not occupy enough excitations when it dominates the exciton transport, causing relatively low $I_{\rm i,o}$. As shown in Fig. 7(b5), adding the cavity drive $p_{\rm c}$ increase largely, but $I_{\rm i,o}$ and $\eta$ do not. Near $\zeta=0$, because of the transfer of excitations the occupation number $p_{11}$ of site 11 increases [Fig. 7(b1)], causing the decrease of $p_{\rm c}$ for HE ST, HE TT, HO-HE ST, and HO-HE TT [Fig. 7(b5)], suppressing the exciton transport of the main channel with the porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) and Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for HO-HE ST and near $\zeta=0.35$ for HO ST. Therefore, the exciton currents $I_{\rm i,o}$ and transport efficiency $\eta$ can be adjusted by the cavity via photons. According to different configurations, we have the conclusion that the coupling type ST is more favorable to exciton transport than TT, and the optimal locations of exciton transport for HO and HE are different in parameter space of $\zeta$, which are consistent with the spectra. In a word, the competition between cavity photons and excitons is the key affecting exciton transport.
@@
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations HE ST and HO-HE ST of $N=9$ [Fig. 8(a3,4)]. The root cause is that the excitation can not enter to multichains through the site 11 [Fig. 8(a1,3)]. It can be seen that for the configurations HE ST and HO-HE ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing with Fig. 6(b3-5). Particularly, at $\delta=0.26$ for HE ST and $\delta=0.36$ for HO-HE ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the crossings of the dressed near zero-energy modes marked by the vertical fine lines in Fig. 4, respectively. The reason for the former is that the excitation on the first and the last column TLSs is transferred to cavity through coherent exciting at the dressed crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ of exciton on the last column TLSs is too few, limiting the exciton transport. At $\delta>0.26$ for HE ST and $\delta>0.36$ for HO-HE ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ on bulk sites [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by the cavity comparing with Fig. 6(b3,4). However, for HO-HE TT both the multichains and the polariton dominate the exciton transport because of the increase of $p_{\rm bulk}$ and $p_{\rm c}$ with $\delta$. For the odd chains, in about $\delta>0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for HE ST and HO-HE ST of $N=8$. The increased $p_{\rm c}$ [Fig. 8(b5)] induces the increase of polaritons, enhancing the exciton transport [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for HE ST of even chain. On the contrary, because of the high $p_{\rm c}$ and $p_{\rm bulk}$ the exciton transport is heightened for HO-HE ST and HO-HE TT of odd chain. Nevertheless, the high $p_{\rm bulk}$ and $p_{\rm c}$ make that the exciton currents and transport efficiency still are high in $\delta>0.7$ for HO-HE ST compared with that under TLS drive. Interestingly, the exciton currents and transport efficiency are significantly large due to the cavity drive in $\delta<0$ for all configurations of $N=9$, which break the universal situations that the exciton currents and transport efficiency are higher in $\delta>0$ than that in $\delta<0$. Specially, when the cavity is driven, the occupation number $p_{11}$ of site 11 is no longer necessary to transport efficiently exciton [Fig. 8(b1)]. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] illustrate that both the multichains and the polaritons participate the exciton transport. Although $p_{11}$ no longer is the condition restricting exciton transport under the TLS+cavity drive, $p_{N}$ still is. Therefore, the low $p_{N}$ limits the exciton transport for HE TT and HO-HE TT of $N=9$ in $\delta<0$. To sum up, in $\delta>0$ the cavity have significant influence to exciton transport, which the exciton transport is suppressed in weak dimerization strength and is enhanced in strong dimerization strength for even chain under TLS drive. In the space of $\delta$, the coupling type ST on even chain is more beneficial to exciton transport. However, when the cavity drive is exerted, ST on odd chain is good for exciton transport.
@@
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 - 5, the structures of exciton transport efficiency are similar to those of spectra. In Fig. 9 the zones of $\zeta$ and $\Delta$ that cavity has obviously effect on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effect these configurations can be classified into two categories. One is HO ST and HO TT that the exciton transport is suppressed, the other is the remaining configurations that the exciton transport is enhanced. For the former the transfer of excitation to the cavity leads to the exciton transport channel with the multichains suppressed [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) $p_{\rm bulk}$ is too small in the zone existing exciton transport in the absence of the cavity, thus the polaritons dominate the transport of exciton, in which the occupation number $p_{\rm c}$ of the cavity photons is relatively large [Fig. 9(ic-lc)]. Further, we can illustrate the mechanism through the spectra. Interestingly, the structures of $p_{\rm c}$ [Fig. 9(gc-lc)] are similar to those of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by the cavity there are the cavity-dressed anticrossings [Fig. 2 and 3], thus LZ transition is the foundation that the cavity effects exciton transport. Particularly, in the high efficient zone of exciton transport there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the dressed anticrossings for HE ST. When the cavity drive is exerted further, for the two categories of configurations the exciton transport is further suppressed and enhanced by the cavity respectively. For HE ST the effect of cavity drive on exciton transport is the biggest than the other configurations, in which the high efficient zone is significantly extended. In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, the effect of the cavity is suppressed for HO-HE TT [Fig. 10(s-v)] because the HO chain and the diagonal inter-chain coupling suppress the dimerized effect. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When the cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiency is strengthened in the zone of the fully dimerized limit and near the zero detuning for HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)]. If the concern is only on the optimal efficiency of exciton transport, HE ST is the preferred configuration in the parameter space of $\zeta$ and $\Delta$ under the TLS drive in the presence of the cavity, as well as HE TT is the preferred configuration in the parameter space of $\delta$ and $\Delta$ under the TLS+cavity drive. Significantly, according to the results the exciton transport can be controlled by the coupling configurations, the cavity, the drive methods, the inter-chain coupling, the dimerization parameter, and the TLS transition frequency detuning.

### v4

@@ IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, the number $N$ and length $L$ of the chains on the exciton transport dynamics by numerically solving the master equation Eq. (13) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by the cavity for strong exciton-cavity coupling, generating the polaritons. To ensure the polariton-dominated exciton transport, we set that the Rabi splitting ($\Omega_{R,k}$) is far stronger than both the cavity ($\kappa$) and TLSs ($\gamma$) decay rates, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
@@ IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relationship between the cavity-dressed energy level crossings (anticrossings) and the exciton transport dynamics to elucidate the underlying transport mechanism. Based on the relation and the properties of exciton transport dynamics, we identify the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations for enhanced exciton transport. Meanwhile, there are two physical quantities to be introduced, the steady-state occupation number $p_{11}$ of site (1,1) and $p_{\rm bulk}$ of bulk sites, where the site (1,1) denotes the first row and column in the TLS array shown in Fig. 1, and the ”bulk” site contains collective occupation of interior sites (columns 2 to N-1). This notation convention applies throughout unless specified otherwise. In addition, according to the Eq. (23) the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ at terminal sites. Hence, subsequent figures display only $I_{\rm o}$ as the transport metric, and $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$.
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. At first glance, one can discover that the inter-chain coupling and dimerization parameter exhibit distinct influences on exciton transport. The $p_{11}$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
@@
For the six configurations, the descending sort of efficiency is successively HO-HE ST, HO ST, HO TT, HO-HE TT, HE ST, HE TT in $\zeta\in[-1,1]$ [Fig. 6(a5)]. For the HO-HE ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess simultaneously relatively high occupation numbers of exciton, which support exciton transport in multichains. Consequently, this behaves the largest exciton transport efficiency, input and outgoing currents [Fig. 6(a3,4,5)]. On the contrary, at $\zeta=0$ and $\zeta\rightarrow\pm 1$, despite more excitations at the site (1,1) [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading to the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [see cyan line in Fig. 6(a2)], which is beneficial to energy storage and has potential applications for designing quantum battery quach2022superabsorption . Except for the HE ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as HO TT and HO-HE TT do near $\zeta=\pm 0.4$. However, for HO ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the abilities both of energy transport and storage are strong. Near $\zeta=\pm 0.7$ for HO ST, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
@@
The system has four configurations HE ST, HE TT, HO-HE ST, HO-HE TT containing HE chain. Because of the different spectra between odd- and even-length HE chain, Fig. 6(b1-5) compares the steady-state dynamics with $\delta$ for $N=8,~{}9$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations HE ST and HO-HE ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the HE TT of $N=9$ is the most beneficial to energy storage because of the high $p_{\rm bulk}$ [Fig. 6(b2)] as well as the lowest $I_{\rm i,o}$ and $\eta$. Particularly, the exciton transport is anomalous in $\delta<0$, where the efficiency is high, but the currents approach to zero, as shown in Fig. 6(b3-5), thus there is no real exciton transport. The reason is that the drive fails to coherently excite the site (1,1), namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the HO-HE ST and HO-HE TT of $N=8,9$ or approaches to zero for the others.
@@
Considering a cavity, we plot the Fig. 7 and 8 corresponding to Fig. 6, which present the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ on the exciton transport dynamics. Our analysis specifically focuses on the change of the steady-state dynamics with $\zeta$ and $\delta$ at the cavity-dressed nearly-zero-energy crossings and anticrossings. In the multichain-cavity coupled system, the main channels of exciton transport contain the multichains and polaritons which depend on occupation number of excitons and photons, while it only is the multichains in the absence of the cavity.
@@
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system in a cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6), respectively. Particularly, to analyse what role the cavity acts as in exciton transport, the dependence of the photon occupation number $p_{\rm c}$ on $\zeta$ is shown in Fig. 7(a5) and 7(b5), in which the extreme points of $p_{\rm c}$ are marked by the vertical fine lines. It can be found through comparison that the extreme points correspond to the dressed nearly-zero-energy crossings and anticrossings in spectra, as well as the discontinuity and extreme points of the second order partial derivative $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of the photon Hopfield coefficients marked by the vertical fine lines in Fig. 2 and 3. In consequence, $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ can track the change of $p_{\rm c}$ with $\zeta$. Comparing with Fig. 6, one can discover that the presence of the cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations HO ST, HE ST, HE TT, and HO-HE ST [Fig. 7(a3,4)]. Moreover, the exciton transport is suppressed by the cavity in the range of $\zeta\neq 0$. For example, $p_{\rm c}$ is high near $\zeta=-0.3$ [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for HO ST [Fig. 7(a3,4)]. On the other hand, polariton-assisted transport boosts $I_{\rm i,o}$ and $\eta$ near $\zeta=0$ for configurations HE ST, HE TT, and HO-HE TT, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number $p_{N}$ on the last column sites [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Because of few excitations injected [Fig. 7(a1)] for HO ST, HO TT and few $p_{N}$ [Fig. 7(a4)] for HO-HE ST near $\zeta=0$, the cavity has no effect on exciton transport. What’s more, it can be seen obviously that the peaks of $p_{\rm c}$ and of $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ the transfer of excitation in the multichains to the cavity suppresses the exciton transport that the multichains server as dominating channel, then causing relatively low $I_{\rm i,o}$. Adding the cavity drive, $p_{\rm c}$ increase dramatically but without proportional $I_{\rm i,o}$ and $\eta$ enhancement as shown in Fig. 7(b5). Because of the transfer of excitations, $p_{11}$ increases near $\zeta=0$ [Fig. 7(b1)], causing $p_{\rm c}$ decrease for HE ST, HE TT, HO-HE ST, and HO-HE TT [Fig. 7(b5)], and suppressing the exciton transport of the main channel with the porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) and Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for HO-HE ST and near $\zeta=0.35$ for HO ST. Therefore, the exciton transport can be adjusted by the cavity photon. For different configurations, we have the conclusion that ST coupling outperforms TT for the exciton transport. Besides, the optimal $\zeta$ ranges differ for HO and HE chains, which are consistent with the spectral features. In a word, the photon-exciton competition governs the suppression and enhancement of exciton transport via cavity.
@@
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity, respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations HE ST and HO-HE ST of $N=9$ [Fig. 8(a3,4)]. The root cause is the blocked excitation injection at the site (1,1) [Fig. 8(a1,3)]. It can be seen that for the configurations HE ST and HO-HE ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing with Fig. 6(b3-5). Particularly, at $\delta=0.26$ for HE ST and $\delta=0.36$ for HO-HE ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the dressed nearly-zero-energy crossings [Fig. 4], respectively. The reason for the former is that the excitations on the edge-column TLSs are transferred to cavity through coherent exciting at the crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ is too few, limiting the exciton transport. At $\delta>0.26$ for HE ST and $\delta>0.36$ for HO-HE ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by the cavity comparing with Fig. 6(b3,4). However, for even-length HO-HE TT both the multichains and the polariton dominate the exciton transport because of the large $p_{\rm bulk}$ and $p_{\rm c}$. For odd-length HO-HE TT, in $\delta\gtrsim 0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for HE ST and HO-HE ST of $N=8$. The increased $p_{\rm c}$ [Fig. 8(b5)] leads to more polaritons, enhancing the polariton channel [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for even-length HE ST. On the contrary, because of the large $p_{\rm c}$ and $p_{\rm bulk}$ the exciton transport is heightened for HO-HE ST and HO-HE TT with odd length, as well as for HO-HE ST in $\delta>0.7$ compared with that under TLS drive. Particularly, the exciton transport currents and efficiency are significantly large due to the cavity drive in $\delta<0$ for all configurations of $N=9$, which break the conventional preference that the transport are superior in $\delta>0$ than that in $\delta<0$. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] demonstrate that both the multichains and the polaritons participate the exciton transport. Interestingly, in the presence of the cavity drive $p_{11}$ is no longer necessary to transport efficiently exciton [Fig. 8(b1)], but $p_{N}$ still is. For instance, the low $p_{N}$ limits the exciton transport for HE TT and HO-HE TT of $N=9$ in $\delta<0$. To sum up, in $\delta>0$ the cavity have significant influence on exciton transport for even-length multichains under TLS drive, which is suppressed in weak $\delta$ and is enhanced in strong $\delta$. In the space of $\delta$, even-length ST is more beneficial to exciton transport than TT under TLS drive, but odd-length ST is under TLS+cavity drive. In addition, no matter which one is main for multichain and polariton channels, the occupation of terminal sites decides the exciton transport.
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 - 5, the effects of $\zeta$, $\delta$, and $\Delta$ on $\eta$ mirror spectral structures.
@@
In Fig. 9 the zones of $\zeta$ and $\Delta$ that the cavity has obviously effects on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effects these coupling configurations of the multichains can be classified into two categories. One is HO ST and HO TT that the exciton transport is suppressed, the others has opposite influence. For the former, the excitation on whole multichains transfer to the cavity disrupts the multichains transport serving as dominating channel [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) too small $p_{\rm bulk}$ limits the multichain channel in the marked zone, but the large $p_{\rm c}$ strengthens the polariton channel [Fig. 9(ic-lc)]. We can further illustrate the mechanism through the spectra. Interestingly, the structures of $p_{\rm c}$ [Fig. 9(gc-lc)] are similar to those of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by the cavity the cavity-dressed anticrossings exist [Fig. 2 and 3], thus LZ transition leads to that the excitations on multichains generate polaritons. Particularly, in the zone with high efficiency there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the cavity-dressed anticrossings for HE ST. When the cavity drive is exerted, for the two categories of configurations the exciton transport is further suppressed and enhanced by the cavity respectively. HE ST shows the biggest effect of cavity drive on exciton transport in six configurations, and that the high efficient zone is significantly extended.
In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, it is contrary for HO-HE TT [Fig. 10(s-v)] because the HO chain and the diagonal inter-chain coupling suppress the dimerized effect. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When the cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiency is strengthened in the zone of $\delta\rightarrow\pm 1$ and near $\Delta=0$ for HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)]. In point of the optimal exciton transport, HE ST is the preferred configuration in the parameter space of $\zeta$ and $\Delta$ under the TLS drive in the presence of the cavity, as well as HE TT is the preferred configuration in the parameter space of $\delta$ and $\Delta$ under the TLS+cavity drive. Significantly, according to the results the exciton transport can be controlled by the coupling configurations, the cavity, the drive methods, the inter-chain coupling, the dimerization parameter, and the TLS transition frequency detuning.